Directions: Answer the following questions on a separate document. Explain how you reached the answer or show your work if a mathematical calculation is needed, or both. Submit your assignment using the assignment link above. This homework assignment is worth 20 points. Assume that as of today, the annualized interest rate on a three-year security is 8 percent, while the annualized interest rate on a two-year security is 6 percent. Use only this information to estimate the one year forward rate two years from now. You need to choose between investing in a a one-year municipal bond with a 6 percent yield and a one year corporate bond with a 9 percent yield. If your marginal federal income tax rate is 20 percent and no other differences exist between these two securities, which one would you invest in? Assume that interest rates for one-year securities are expected to be 3 percent today, 6 percent one year from now, and 7 percent two years from now. Using only the pure expectations theory, what are the current interest rates on two-year and three-year securities? Identify some factors that influence the shape of the yield curve. Describe how financial market participants use the yield curve to make decisions. Assume there is a sudden expectation of increasing interest rates in the future. What would be the effect on the shape of the interest curve? Explain. Identify the primary objectives of the Federal Open Market Committee and the means by which it attempts to achieve these objectives. What is the Beige Book and why do market participants pay attention to it? Is it the role of the Fed or Congress to determine the fate of large financial institutions that are near bankruptcy? Does the Federal Reserve directly or indirectly influence equity security prices? Based on your response, do equity market participants focus on current Fed actions or the expected ones when it comes to pricing equity securities? Visit the Consumer Financial Protection Bureau's website. Identify one policy objective and its effects on consumers and businesses.
Answer #1

A.One year fwd rate 2 years from now = (1.08^3/1.06^2) - 1= 12.1%

B. The interest income earned from most municipal bonds is exempt from taxes

After tax interest in corporate bond = 9*(1-0.2)= 7.2%

Since after tax interest on corporate bond is higher than that on a municipal bond, it is preferable to invest in the corporate bond.

C.

D. Shape of yield curve depends on the following factors:

1. Supply and demand

2. Liquidity premium expectation

3. Interest rate expectation

4. Changes in economic output and growth

E. A normal yield curve has higher yield on longer maturity bonds due to inherent risk with time, and this happens in a normal economy. in case of recession, the yield curve gets inverted and lower yields are expected on longer maturity bonds.